The Medal Plus Company manufactures windows. Its manufacturing plant has the capacity to produce 20,000 windows each month. C300,000 150,000 Variable costs that vary with number of units produced Direct materials Direct manufacturing labour VariableMedal Plus has just received a special one-time-only order for 5,000 windows at $125 per window. Accepting the special order1. Should Medal Plus accept this special order? Show your calculations. 2. Suppose plant capacity were only 17,500 windows inRequirement 1. Should Medal Plus accept this special order? Show your calculations. Begin by completing an analysis and startRevenues Variable costs: Direct materials Direct manufacturing labour Batch manufacturing costs Fixed costs: Fixed manufactur
Answer #1

Requirement 1

Particulars Without Accepting Special Order After Accepting Special Order Increase / ( Decrease)
Revenue 22,50,000 28,75,000 6,25,000
Variable Costs
Direct Materials 3,00,000 4,00,000 1,00,000
Direct Labour 1,50,000 2,00,000 50,000
Variable Costs 3,75,000 5,00,000 1,25,000
Total Variable Costs 8,25,000 11,00,000 2,75,000
Contribution Margin 14,25,000 17,75,000 3,50,000
Fixed Costs
Fixed Manufacturing Costs 3,00,000 3,00,000 0
Fixed Marketing Costs 55,000 55,000 0
Total Fixed Costs 3,55,000 3,55,000 0
Operating Income 10,70,000 14,20,000 3,50,000


Yes the Company should accept the Special Order as the operating income is increased by $ 350,000.

Notes and Supporting Work

Contribution Margin = Sales - Total Variable Costs

Operating Income = Contribution Margin - Total Fixed Costs

Calculation for Accepting the Sales Order

Sales Revenue = 15,000 Units * $ 150 = $ 22,50,000

Sales Revenue for additional sales of 5,000 Units = 5,000 * $ 125 per Unit = $ 625,000

Total Sales Revenue = $ 28,75,000

Total Variable Costs

Variable Manufacturing Overhead = 400 Batches * $ 1,250 per Batch = $ 600,000

Direct Labour = $ 10 per Unit * 20,000 Units = $ 200,000

Direct Materials = $ 20 per Unit * 20,000 Units = $ 400,000

Additional Information

Direct Materials Cost per Unit = 300,000 / 15,000 = $ 20 per Unit

Direct Labour Cost per Unit = 150,000 / 15,000 = $ 10 per Unit

Calculation for 15,000 Units without special order has already been done in the question.

Increase / (Decrease) = Values after Acceptance of Special Order - Values without Accepting Special Order

Requirement 2

Particulars    Without Accepting Special Order After Accepting Special Order Increase / ( Decrease)
Revenue 22,50,000 25,00,000 2,50,000
Variable Costs
Direct Materials 3,00,000 3,50,000 50,000
Direct Labour 1,50,000 1,75,000 25,000
Variable Costs 3,75,000 4,37,500 62,500

Total Variable Costs

8,25,000 9,62,500 1,37,500

Contribution Margin

14,25,000 15,37,500 1,12,500
Fixed Costs
Fixed Manufacturing Costs 3,00,000 3,00,000 0
Fixed Marketing Costs 55,000 55,000 0
Total Fixed Costs 355,000 355,000 0
Operating Income 10,70,000 11,82,500

1,12,500

Yes the Company should accept the Special Order as the operating income is increased by $ 112,500

Notes and Supporting Work

Contribution Margin = Sales - Total Variable Costs

Operating Income = Contribution Margin - Total Fixed Costs

Calculation for Accepting the Sales Order

Accepting Special Order will reduce the external sales by 2,500 Units

Sales Revenue = 12,500 Units * $ 150 = $ 18,75,000

Sales Revenue for additional sales of 5,000 Units = 5,000 * $ 125 per Unit = $ 625,000

Total Sales Revenue = $ 25,00,000

Total Variable Costs

Variable Manufacturing Overhead = 4)350 Batches * $ 1,250 per Batch = $ 437,500

Direct Labour = $ 10 per Unit * ,17,500 Units = $ 175,000

Direct Materials = $ 20 per Unit * 17,500 Units = $ 350,000

Additional Information

Direct Materials Cost per Unit = 300,000 / 15,000

Direct Labour Cost per Unit = 150,000 / 15,000 = $ 10 per Unit

Calculation for 15,000 Units without special order has already been done in the question.

Increase / (Decrease) = Values after Acceptance of Special Order - Values without Accepting Special Order

Requirement 3

Particulars Without Accepting Special Order After Accepting Special Order Increase / ( Decrease)
Revenue 22,50,000 26,50,000 4,00,000
Variable Costs
Direct Materials 3,00,000 4,00,000 1,00,000
Direct Labour 1,50,000 2,00,000 50,000
Variable Costs 3,75,000 5,00,000 1,25,000
Total Variable Costs 8,25,000 11,00,000 2,75,000
Contribution Margin 14,25,000 15,50,000 1,25,000
Fixed Costs
Fixed Manufacturing Costs 300,000 3,00,000 0
Fixed Marketing Costs 55,000 55,000 0
Total Fixed Costs 355,000 3,55,000 0
Operating Income 10,70,000 11,95,000

1,25,000

Yes the Company should accept the Special Order as the operating income is increased by $ 125,000.

Notes and Supporting Work

Contribution Margin = Sales - Total Variable Costs

Operating Income = Contribution Margin - Total Fixed Costs

Calculation for Accepting the Sales Order

Accepting Special Order will reduce the External Unit Sales Price by $ 15

Sales Revenue = 15,000 Units * $ 135 = $ 20,25,000

Sales Revenue for additional sales of 5,000 Units = 5,000 * $ 125 per Unit = $ 625,000

Total Sales Revenue = $ 26,50,000

Total Variable Costs

Variable Manufacturing Overhead = 400 Batches * $ 1,250 per Batch = $ 500,000

Direct Labour = $ 10 per Unit * 20,000 Units = $ 200,000

Direct Materials = $ 20 per Unit * 20,000 Units = $ 400,000

Additional Information

Direct Materials Cost per Unit = 300,000 / 15,000 = $ 20

Direct Labour Cost per Unit = 150,000 / 15,000 = $ 10

Increase / (Decrease) = Values after Acceptance of Special Order - Values without Accepting Special Order